OpenStudy (anonymous):

find all points P on the parabola y=x^2 such that the tangent line at P passes through the point (0,-4).

OpenStudy (anonymous):

I've got it never mind.

OpenStudy (anonymous):

\[ \ \ \ y(x)=x^2\\ \text{First, differentiate with respect to }x.\\ \ \ \ y'(x)=2x\\ \text{We know that the point-slope form of a tangent}\\ \text{line equation is as follows.}\\ \ \ \ y-y_0=y'(x_0)(x-x_0)\\ \ \ \ y-y_0=2x_0(x-x_0)\\ \ \ \ y-y_0=2x_0x-2x_0^2\\ \ \ \ y-x_0^2=2x_0x-2x_0^2\\ \ \ \ y=2x_0x-x_0^2\\ \text{We know that our lines must pass through}\\ \ \ \ (x,y)=(0,-4)\\ \text{So we may reduce the tangent line equation as follows.}\\ \ \ \ -4=2(0)x_0-x_0^2\\ \ \ \ 4=x_0^2\\ \ \ \ x_0=\pm2\\ \text{Now that we have the solutions for }x_0\text{ we can find }y_0.\\ \ \ \ y_0 = y(\pm2)=(\pm2)^2=4\\ \text{Therefore our solution is...}\\ \ \ \ (x_0,y_0)\in\{(-2,4),(2,4)\} \]